[figure: 9.7 cm]

TEMPERATURE DEPENDENCE OF NMR IN MAGNETITE

P. Novák^{a}, J. Englich^{b}, J. Kohout^{b} and H. Štěpánková^{b} aInstitute of Physics, Academy of Sciences of the Czech Republic Cukrovarnická 10, 162 53 Praha 6, Czech Republic bFaculty of Mathematics and Physics, Charles University, V Holešovičkách 2, 180 40 Praha 8, Czech Republic

We report the temperature dependence of ^{57}Fe NMR in a single crystal of Fe_{3}O_{4} measured by the spin echo method between 4.2 and 140 K (temperature of the Vervey transition T_{V} = 121 K). We were able to register all lines detected at 4.2 K also at higher temperatures up to T_{V}. The spectra are compatible with the space group Cc, which predicts 8 and 16 lines for iron on tetrahedral (A) and octahedral (B) sites, respectively. One of the Fe(A) lines is doubly degenerate for all T ≤ T_{V}, which fact restricts the possible models of the Fe^{2+} and Fe^{3+} ordering on the B sites. No traces were found of the motionally narrowed spectrum below T_{V}, which would correspond to the electrons on the B sites moving faster comparing to the characteristic time of the NMR measurement (10^{−8} s). At T_{V} the spectrum is simplified to two lines, one corresponding to Fe(A), the other to Fe(B) as expected for magnetization parallel to ⟨001⟩. Above T_{V} no remnant of the low temperature spectrum was found, indicating that if there is a short range order, it fluctuates faster than the NMR frequency.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : P. Novák

Address for correspondence : Institute of Physics Cukrovarnická 10, 162 53 Praha 6 Czech Republic

Email address : novakp@fzu.cz